The escape speed from a distant planet is 45 km/s. If the acceleration of gravity on the planet surface is 14.0 m/s2, what is the radius of the planet?
Answer #1

given

ve = 45 km/s = 45*10^3 m/s

g = 14 m/s^2

let R is the radius of the planet.

we know escape speed on a planet,

ve = sqrt(2*g*R)

ve^2 = 2*g*R

==> R = Ve^2/(2*g)

= (45*10^3)^2/(2*14)

= 7.23*10^7 m